OpenStudy (anonymous):

Which 2 consecutive integers have a sum of -105

OpenStudy (pfenn1):

define n=an integer, the next consecutive integer would be n+1 so n + (n+1) = -105 solve for n

OpenStudy (anonymous):

I dont know if i knew would i ask the question??!

OpenStudy (pfenn1):

you don't know how to solve for n? n + (n+1) = 2n + 1=-105 subtract 1 from each side of the equation 2n = -106 divide each side by 2 n = -53 n+1 = -52 so the two consecutive integers are -53 and -52 does that help?